OpenStudy (anonymous):

Evaluate the expression below when x = 3. 5x^2-12 / x+3 a. 3 b. 11 / 3 c. 11 / 2 d. 26 / 3 please help me.

OpenStudy (anonymous):

3 squared = 9 times 5 = 45 minus 12 = 33. 3+3 =6. so, it's 33/6. here

OpenStudy (anonymous):

there is no 33/6 o.o lol

OpenStudy (aric200):

:O

OpenStudy (anonymous):

you guys got anything? :C

OpenStudy (aric200):

5(3)^2-12/3+3

OpenStudy (aric200):

hold on

OpenStudy (anonymous):

alright

jimthompson5910 (jim_thompson5910):

alliecat123, you need to reduce that fraction

OpenStudy (aric200):

Ah

OpenStudy (anonymous):

its B then?

OpenStudy (aric200):

?

OpenStudy (aric200):

Facepalm

OpenStudy (anonymous):

is it B?

OpenStudy (aric200):

Divide both by 3

OpenStudy (aric200):

33/3 = 11 6/3=2

OpenStudy (aric200):

C.

OpenStudy (anonymous):

thanks :)

OpenStudy (aric200):

I did nothing

OpenStudy (aric200):

Just reduced

OpenStudy (anonymous):

k i gave her the medal for helping :)

OpenStudy (aric200):

Allie should get medal :D

OpenStudy (anonymous):

i did :)

OpenStudy (aric200):

It hurt getting a medal then losing it :O

OpenStudy (anonymous):

xd i fanned you :P

OpenStudy (aric200):

Thank you it is a little warm in here.